OpenStudy (anonymous):

one number exceeds another by 20. The sum of the number is 66. What are the numbers?

OpenStudy (anonymous):

let the numbers be A and B, so say A exceeds B by 20. Then instead of saying "A", you could say "B+20" Would you agree?

OpenStudy (anonymous):

Since we don't know the numbers let x and y be the numbers with x > y.

OpenStudy (anonymous):

So, the sum of A and B is 60....... 60 = A + B But you can call "A" as "B+20" so 60 = (B+20) + B

OpenStudy (anonymous):

oops, meant to write 66, not 60

OpenStudy (anonymous):

Then since x > y, it implies x - y = 20

OpenStudy (anonymous):

66 = (b+20) + b 66 = 2b + 20 46 = 2b b = 23 a = b+20, so a = 23+20 = 43 The two numbers are 43 and 23 (Check: 23 + 43 = 66)

OpenStudy (anonymous):

Also x + y = 66 This gives us two equations in two variables Hence, solve x - y = 20 and x + y = 66 simultaneously

OpenStudy (anonymous):

If this is early algebra, they may not be at 2 variable problems yet. Often they seem to teach just direct substitution based on the relationship without showing the rigor of the system of 2 equations. But I wish they'd teach it the way you are :)

OpenStudy (anonymous):

You can also use the one variable state approach

OpenStudy (anonymous):

Thanks JakeV8 for the appreciation.